What is the value of the following Python expression?

2 * 1 ** 2 ** 1 ** 2?
Answer #1
2 * 1 ** 2 ** 1 ** 2
= 2 * 1 ** 1 ** 2
= 2 * 1 ** 2
= 2 * 1 
= 2

Answer

2